Problem 16-15 MM and Taxes [LO2] Meyer & Co. expects its EBIT to be $123,000 every year forever. The firm can borrow at 7 percent. The company currently has no debt, and its cost of equity is 13 percent and the tax rate is 25 percent. The company borrows $174,000 and uses the proceeds to repurchase shares. a. What is the cost of equity after recapitalization? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1


Value unlevered =EBIT*(1-Tax Rate)/Cost of Equity =123000*(1-25%)/13% =709615.3846
Value of firm levered =Value Unlevered +Debt*Tax Rate = 709615.3846+174000*25% =753115.3846
New Equity =753115.3846-174000 =579115.3846
a. Cost of equity after capitalization =Cost of Equity+(Cost of Equity -Cost of Debt)*(Debt/Equity)*(1-Tax Rate)
=13%+(13%-7%)*174000/579115.3846 *(1-25%) =14.3521% or 14.35%

b.WACC =Weight of Equity*Cost of Equity+Weight of Debt*Cost of Debt*(1-Tax Rate)
=579115.3846/753115.3846*14.3521%+17400/753115.3846*7%*(1-25%) =11.16%